OpenStudy (anonymous):

Rewrite the following expressions in x=my + b form. a. 3x + 6y + 9 = 0 b. -2x + 6y - 4 = 0 c. -4x - 2y = 0

OpenStudy (anonymous):

Set everything equal to x, then you'll get it to x=my + b

OpenStudy (anonymous):

umm..I don't get what you mean! sorry

OpenStudy (anonymous):

I'll start the first one for you

OpenStudy (anonymous):

a. Subtract 9 from both sides so that you get: 3x+6y= -9 Then subtract 6y from both sides to get: 3x= -9-6y Then divide by 3 on both sides to get: x= (-9-6y)/3 Since 9 and 6 can be divided by 3, x= -3-2y

OpenStudy (anonymous):

to move everything from the left side to the right side, you need to either add/subtract from both sides

OpenStudy (anonymous):

since you need it in x= my+b, (a) should look like this: x = -2y-3